# Evaluating and Steering Modality Preferences in Multimodal Large Language Model

### v2

@@ 6 Conclusion
This paper investigates modality preference in multimodal large language models (MLLMs). We carefully curate a modality conflict dataset and use a controlled experimental setup to quantitatively evaluate modality preference. Besides, we find the direction of modality preference can be captured within the latent representations of MLLMs. Inspired by this, we propose a modality preference probing and steering method, which enables significant and flexible changes in modality preference. Experiments show that the proposed method generalizes well to downstream tasks, such as multimodal machine translation and multimodal understanding tasks.
Appendices
All codes, data, and instructions for our MC2 can be found in https://anonymous.4open.science/r/Modality-Preference-8016. MC2 is released under a Creative Commons Attribution 4.0 License (CC BY 4.0).
Our supplementary materials are summarized as follows:
@@ A.1 Limitations
This paper investigates the modality preference in multimodal large language models (MLLMs) using a controlled experiment setup with a modality conflict dataset. In constructing the dataset, we employs LLaVA1.5-7B and QwenVL-7B to filter samples and ensure that most models could answer questions correctly based on a single modality. However, this process requires multiple iterations and turned out to be time-consuming. Therefore, devising a more efficient and elegant method for sample selection may be of greater importance.
A.2 Social Impacts
@@ A.3 Use of LLM
In this work, we use the LLMs including GPT-4o-mini (Hurst et al., 2024a) and DeepSeekV3 (Liu et al., 2024a), and MLLMs including LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023) to help annotate the text context in MC2, as detailed in Section B. We evaluate the modality preference of 20 open-source MLLMs and GPT-4o-mini (Hurst et al., 2024a), and steer the modality preference of Qwen2VL-7B (Wang et al., 2024), Qwen2.5VL-7B (Bai et al., 2025), and LLaVA-OneVision-7B (Li et al., 2024a) and InternVL3-8B (Zhe et al., 2024). Besides, we also utilize the LLMs to correct the grammatical errors.
@@ B.1 Conflict Text Context Generation
Conflict Context Generation for counting task using DeepSeekV3 Instruction: # Given a description of an image and a corresponding counting type question with its answer, now you are required to generate a text context that points to an answer that fluctuates by 1 or 2 from the original answer. The context explicitly supports the new answer, providing clear evidence that aligns logically with the counting question. Only one alternative answer should be generated. Caption: {caption} Question: {question} Answer: {answer based on vision context} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
Conflict Context Generation of for other tasks using DeepSeekV3 Instruction: # Given the caption of an image and a corresponding {task-type} type question with its answer, now you are required to generate a text context as a premise that supports a new distractor answer for the question. The context should mimic the environment described in the caption but should not include {answer based on vision context}, while maintaining logical consistency within the context. Only one alternative answer should be generated. Caption: {caption} Question: {question} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
Conflict Context Generation for other tasks using GPT-4o-mini Instruction: # Given a caption of an image and a corresponding counting question with its answer, you are required to generate a single text context that provides an indirect premise leading to a new answer that fluctuates by 1 or 2 from the original answer. The context should build an indirect premise to the new answer. Carefully design this context. For this task, I want you to first describe the scene with a certain quantity and then introduce an increase or decrease in that quantity to imply the final answer and don’t include the final answer. Only one alternative answer should be generated. Caption: {caption} Question: {question} Answer: {answer based on vision context} Task-type: {task-type} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
Conflict Context Generation for count task using GPT-4o-mini Instruction: # Given the caption of an image and a corresponding question with its answer, now you are required to generate a text context as the indirect premise of a new answer for the question, which belongs to the same category as the original answer. The context should support the new answer, include the caption while maintaining logical consistency within the context and don’t include the final answer. Only one alternative answer should be generated. Caption: {caption} Question: {question} Answer: {answer} Task-type: {task-type} Output the new answer enclosed in <answer> </answer> and the context enclosed in <context> </context> tags.
@@
  • Modality-Answer Alignment. First, for each context from different modalities (image and text), annotators assess whether it independently provides sufficient information to correctly answer the question. This step is particularly important because the original VQA answers in the TDIUC (Kafle & Kanan, 2017) dataset may contain error annotations, and the LLM-generated contexts and answers may occasionally be inconsistent.
@@ C.1 Evaluatiion Detail for Modality Preference
We assess open-source multimodal large language models (MLLMs) with different parameter sizes, including LLaVA1.5-7B/13B (Liu et al., 2024b), LLama3.2-11B-Vision-Instruct (Grattafiori et al., 2024), LLaVA-OneVision-7B/72B (Li et al., 2024a), CogVLM2-19B (Hong et al., 2024a), mPLUG-Owl3-24-07 (Ye et al., 2024), Qwen2VL-7B (Wang et al., 2024), GLM-4V-9B (Du et al., 2022), SPHINX-V2-1K (Lin et al., 2023), InternVL3-9B/14B/38B/78B (Zhe et al., 2024), LLaVA-next-7B/13B/34B (Liu et al., 2024b) and Qwen2.5VL-7B/32B/72B (Bai et al., 2025). All the open-source models are evaluated using NVIDIA A100 or A800 GPUs. We also evaluate the proprietary model, GPT-4o-mini (Hurst et al., 2024a) via the official API.
@@ C.2 Can the Vision Ratio provide guidance for downstream task performance?
We evluate the performance of Qwen2.5VL-7B, Qwen2.5VL-72B, InternVL3-9B, InternVL3-14B, InternVL3-38B, InternVL3-78B, LLaVA-OneVison-7B, LLaVA-OneVison-72B, LLava-next-7B, LLava-next-13B on 7 general multimodal understanding benchmarks including MMMU (Yue et al., 2024), MME (Chaoyou et al., 2023), MMBench (Liu et al., 2024d), RealwordQA (X.AI, 2024), MMStar (Masry et al., 2022), InfoVQA (X.AI, 2024) and ChartQA (Masry et al., 2022). We compute the average score on all datasets, where MME score is normalized between 0-1, as shown in Table 5.
@@ C.3 The Details for Controlling Modality Preference
In the left panel of Figure 4, we provide the Vision Ratio results for LLaVA-OneVision-7B, Qwen2.5VL-7B, Qwen2VL-7B and InternVL3-8B. We also present more results on controlling modality preference through instruction design for preference towards the vision modality and the text modality in Table 7 and Table 8. For each setting, we report the results measured by $S_{vision}$, vision-based accuracy and $S_{text}$, text-based accuracy.
@@ D.1 Details for Pattern of hidden states
In the main text, we visualize the layer-wise absolute difference and standard deviation of the hidden states for Qwen2.5VL-7B. As shown in Figure 10, we present the visualization of hidden states for LLaVA-OneVision-7B, Qwen2VL-7B, and InternVL3-8B. For each model, we selected layers with large absolute differences and small standard deviations. This means we identified the layers that showed stable and significant differences between instructions with modality preference towards vision context and text context, which are then used to steer and adjust the model’s modality preference.
Figure 10: Layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction for LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-8B models from left to right.
D.2 Evaluation of visual understanding and multimodal multimodal machine translation
@@
Ambigcaps (Li et al., 2021) benchmark explores the role of datasets in stimulating the leverage of the visual modality and proposes methods to highlight the importance of visual signals in the datasets. We evaluate the multimodal machine translation task on this dataset using the Qwen2.5VL-7B model, with different instructions that favor the vision modality and the text modality. Conversely, when guided toward the text modality, the model places greater emphasis on the source language, leading to more accurate grounding in multimodal machine translation. This adjustment prevents the model from over-relying on visual content and from introducing spurious objects or extraneous details into the translation output.

### v3

@@ 6 Conclusion
This paper investigates modality preference in multi-modal large language models (MLLMs). We carefully curate a modality conflict dataset and use a controlled experimental setup to quantitatively evaluate modality preference. Besides, we find the direction of modality preference can be captured within the latent representations of MLLMs. Inspired by this, we propose a modality preference probing and steering method, which enables significant and flexible changes in modality preference. Experiments show that the proposed method generalizes well to downstream tasks, such as multi-modal reasoning and understanding tasks.
Impact Statement
@@ A.1 Limitations
This paper investigates the modality preference in multi-modal large language models (MLLMs) using a controlled experiment setup with a modality conflict dataset. In constructing the dataset, we employs LLaVA1.5-7B and QwenVL-7B to filter samples and ensure that most models could answer questions correctly based on a single modality. However, this process requires multiple iterations and turned out to be time-consuming. Therefore, devising a more efficient and elegant method for sample selection may be of greater importance.
A.2 Strategic Insights for Future Research
Based on the experiment analysis, we highlight two key principles for designing effective multi-modal learning strategies: 1) As observed in Section 4.2, increasing the proportion of data with multi-modal context (TVQA) accelerates the adjustment of modality preference. Therefore, for MLLMs—especially those inheriting parameters from LLMs—introducing a more TVQA data can more rapidly mitigate the inherent language bias, facilitating faster convergence toward balanced multi-modal learning.
2) We identify that the Vision Ratio serves as a reliable indicator of visual understanding, and adjusting modality preference improves downstream performance. Consequently, combined with the first point, selecting an appropriate ratio of TVQA training data can adjust modality preferences to satisfy the specific requirements of different downstream.
@@ A.3 Use of LLM
In this work, we use the LLMs including GPT-4o-mini (Hurst et al., 2024a) and DeepSeekV3 (Liu et al., 2024a), and MLLMs including LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023) to help annotate the text context in MC2, as detailed in Section B. We evaluate the modality preference of 20 open-source MLLMs and GPT-4o-mini (Hurst et al., 2024a), and steer the modality preference of Qwen2VL-7B (Wang et al., 2024), Qwen2.5VL-7B (Bai et al., 2025), and OneVision-7B (Li et al., 2024a) and InternVL3-8B (Zhe et al., 2024). Besides, we also utilize the LLMs to correct the grammatical errors.
@@ B.1 Conflict Text Context Generation
Conflict Context Generation for counting task using DeepSeekV3 Instruction: # Given a description of an image and a corresponding counting type question with its answer, now you are required to generate a text context that points to an answer that fluctuates by 1 or 2 from the original answer. The context explicitly supports the new answer, providing clear evidence that aligns logically with the counting question. Only one alternative answer should be generated. Caption: {caption} Question: {question} Answer: {answer based on vision context} Output the new answer enclosed in ¡answer¿ ¡/answer¿ and the context enclosed in ¡context¿ ¡/context¿ tags.
Conflict Context Generation of for other tasks using DeepSeekV3 Instruction: # Given the caption of an image and a corresponding {task-type} type question with its answer, now you are required to generate a text context as a premise that supports a new distractor answer for the question. The context should mimic the environment described in the caption but should not include {answer based on vision context}, while maintaining logical consistency within the context. Only one alternative answer should be generated. Caption: {caption} Question: {question} Output the new answer enclosed in ¡answer¿ ¡/answer¿ and the context enclosed in ¡context¿ ¡/context¿ tags.
Conflict Context Generation for other tasks using GPT-4o-mini Instruction: # Given a caption of an image and a corresponding counting question with its answer, you are required to generate a single text context that provides an indirect premise leading to a new answer that fluctuates by 1 or 2 from the original answer. The context should build an indirect premise to the new answer. Carefully design this context. For this task, I want you to first describe the scene with a certain quantity and then introduce an increase or decrease in that quantity to imply the final answer and don’t include the final answer. Only one alternative answer should be generated. Caption: {caption} Question: {question} Answer: {answer based on vision context} Task-type: {task-type} Output the new answer enclosed in ¡answer¿ ¡/answer¿ and the context enclosed in ¡context¿ ¡/context¿ tags.
Conflict Context Generation for count task using GPT-4o-mini Instruction: # Given the caption of an image and a corresponding question with its answer, now you are required to generate a text context as the indirect premise of a new answer for the question, which belongs to the same category as the original answer. The context should support the new answer, include the caption while maintaining logical consistency within the context and don’t include the final answer. Only one alternative answer should be generated. Caption: {caption} Question: {question} Answer: {answer} Task-type: {task-type} Output the new answer enclosed in ¡answer¿ ¡/answer¿ and the context enclosed in ¡context¿ ¡/context¿ tags.
@@
  • Modality-Answer Alignment. First, for each context from different modalities (image and text), annotators assess whether it independently provides sufficient information to correctly answer the question. This step is particularly important because the original VQA answers in the TDIUC (Kafle and Kanan, 2017) dataset may contain error annotations, and the LLM-generated contexts and answers may occasionally be inconsistent.
@@ C.1 Evaluatiion Detail for Modality Preference
We assess open-source multi-modal large language models (MLLMs) with different parameter sizes, including LLaVA1.5-7B/13B (Liu et al., 2024b), LLama3.2-11B-Vision-Instruct (Grattafiori et al., 2024), OneVision-7B/72B (Li et al., 2024a), CogVLM2-19B (Hong et al., 2024a), mPLUG-Owl3-24-07 (Ye et al., 2024), Qwen2VL-7B (Wang et al., 2024), GLM-4V-9B (Du et al., 2022), SPHINX-V2-1K (Lin et al., 2023), InternVL3-9B/14B/38B/78B (Zhe et al., 2024), LLaVA-next-7B/13B/34B (Liu et al., 2024b) and Qwen2.5VL-7B/32B/72B (Bai et al., 2025). All the open-source models are evaluated using NVIDIA A100 or A800 GPUs. We also evaluate the proprietary model, GPT-4o-mini (Hurst et al., 2024a) via the official API.
@@ C.2 Can the Vision Ratio provide guidance for downstream task performance?
We evluate the performance of Qwen2.5VL-7B, Qwen2.5VL-72B, InternVL3-9B, InternVL3-14B, InternVL3-38B, InternVL3-78B, LLaVA-OneVison-7B, LLaVA-OneVison-72B, LLava-next-7B, LLava-next-13B on 7 general multi-modal understanding benchmarks including MMMU (Yue et al., 2024), MME (Chaoyou et al., 2023), MMBench (Liu et al., 2024d), RealwordQA (X.AI, 2024), MMStar (Masry et al., 2022), InfoVQA (X.AI, 2024) and ChartQA (Masry et al., 2022). We compute the average score on all datasets, where MME score is normalized between 0-1, as shown in Table 5.
@@ C.3 The Details for Controlling Modality Preference
In the left panel of Figure 4, we provide the Vision Ratio results for OneVision-7B, Qwen2.5VL-7B, Qwen2VL-7B and InternVL3-8B. We also present more results on controlling modality preference through instruction design for preference towards the vision modality and the text modality in Table 14 and Table 15. For each setting, we report the results measured by $S_{\text{vision}}$, vision-based accuracy and $S_{\text{text}}$, text-based accuracy.
@@ D.1 Details for Pattern of hidden states
In the main text, we visualize the layer-wise absolute difference and standard deviation of the hidden states for Qwen2.5VL-7B. As shown in Figure 11, we present the visualization of hidden states for OneVision-7B, Qwen2VL-7B, and InternVL3-8B. For each model, we selected layers with large absolute differences and small standard deviations. This means we identified the layers that showed stable and significant differences between instructions with modality preference towards vision context and text context, which are then used to steer and adjust the model’s modality preference.
Figure 11: Layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction for OneVision-7B, Qwen2VL-7B and InternVL3-8B models from left to right.
D.2 Evaluation of visual understanding and multi-modal multi-modal machine translation
@@
Ambigcaps (Li et al., 2021) benchmark explores the role of datasets in stimulating the leverage of the visual modality and proposes methods to highlight the importance of visual signals in the datasets. We evaluate the multi-modal machine translation (MMT) task on this dataset using the Qwen2.5VL-7B model. multi-modal contexts in MMT are both complementary and contradictory: the visual information provides helpful context for translation, but the potential for conflicting, non-visual signals can interfere with grounding the source language. Consequently, the proposed method is designed to steer the modality preference toward the text modality to ensure robustness against these visual-textual conflicts.
@@
Conversely, when guided toward the text modality, the model places greater emphasis on the source language, leading to more accurate grounding in multi-modal machine translation. This adjustment prevents the model from over-relying on visual content and from introducing spurious objects or extraneous details into the translation output.
Appendix E More Experiment Analysis
E.1 The Sensitivity Analysis for the Evaluation of Modality Preference
To verify whether the current 2k-sample scale of $MC^{2}$ is sufficient to ensure the stability of both the preference evaluation and steering results, we conduct a sensitivity analysis by randomly selecting a specified quantity of samples from each category for assessment. As shown in Table 6, we calculate the Vision Ratio for OneVision-7B and Qwen2.5VL-7B. Results demonstrate that as the sample size increases per category, the Vision Ratio begins to stabilize around 150 samples. These experiments suggest that the current dataset size for each task is sufficient to ensure the stability of evaluation of modality preference.
In the future, we would like to expand MC2 with a wider variety of tasks (e.g., texture recognition) and modalities (e.g., Audio), further enhancing its comprehensiveness and generalizability.
Table 6: The sensitivity analysis for the evaluation of modality preference. We  randomly select a specified quantity of samples from each category for assessment, measured by Vision Ratio.
Model 25 50 75 100 125 150 175 200 225 250
OneVision-7B 29.2 28.0 29.1 29.6 28.1 26.4 26.6 26.7 26.7 26.3
Qwen2.5VL-7B 56.7 59.4 60.2 59.9 60.4 59.8 60.2 59.7 59.8 59.6
E.2 More Results for Downstream Tasks
E.2.1 More results for Phd dataset
We evaluate the performance of Phd dataset using the proposed method for LLaVA-1.6-7B and the results are in Table 7. We observe that LLaVA-1.6-7B, due to its severe text preference, only achieves an average ACC score of $0.7$ on the PHD-icc subset. Applying our method significantly boosts the model’s performance across the two subsets, thereby demonstrating the cross-model generalization of our approach.
Table 7: Performance of the proposed method on the visual understanding benchmark, Phd using LLaVA-1.6-7B.
Phd Method Attribute Sentiment Positional Counting Object Avg
Phd-icc LLaVA-1.6-7B 0.5 0.0 0.0 1.5 1.5 0.7
InstDesign 2.0 0.5 0.5 1.5 9.5 2.8
Ours 3.5 1.5 1.0 3.0 15.0 4.8
Phd-iac LLaVA-1.6-7B 5.5 8.0 4.0 10.5 29.0 11.4
InstDesign 7.5 12.5 14.5 15.5 44.5 18.9
Ours 11.5 20.0 20.5 22.0 51.0 25.0
E.3 Ablation Study
In this section, we conduct the detailed ablation study to analyze the proposed method.
E.3.1 The number of probing samples
We compute the preference direction in Equation 1 using varied sample sizes but test the steering performance on the complete $MC^{2}$ dataset. We report the $S_{\text{Vision}}$ and $S_{\text{Text}}$ results for OneVision-7B and Qwen2.5VL-7B in Table 8. We observe that steering performance remains stable even when the steering vector is derived from a limited number of samples.
Table 8: The ablation study for the varied sample number of computing the preference direction.
Model 25 50 75 100 125 150 175 200 225 250
OneVision-7B 56.0 55.9 56.3 57.1 57.0 57.3 57.3 57.4 57.2 57.1
Qwen2.5VL-7B 62.2 62.2 61.7 62.6 64.0 62.8 62.7 63.1 62.8 63.6
E.3.2 The diversity of probing samples
To study the impact of data diversity for probing task, we experiment by using only a single task for probing and applying the resulting vector to steer all other tasks. We report the $S_{\text{Vision}}$ and $S_{\text{Text}}$ for OneVision-7B and Qwen2.5VL-7B for entire dataset in Table 9. We observe that OneVision-7B achieves competitive performance compared to our initial implementation for using nearly each probing task, with Qwen2.5VL-7B showing similar success on over half the tasks. Further analysis finds that the most effective single-probing tasks are those where the initial modality preference change was more pronounced.
Table 9: The ablation study for the varied diversities of computing the preference direction.
Model Sport Attribute Sentiment Positional Counting Color Activity Object Ours
OneVision-7B 58.6 55.8 59.3 57.8 59.0 53.4 58.5 58.6 57.1
Qwen2.5VL-7B 64.1 59.5 70.0 51.9 47.4 65.5 63.0 60.6 63.6
E.3.3 Different steering intensities
To investigate the performance with varied steering intensities, we introduce a scaling coefficient $\lambda$ to the steering weight $w$ in Equation 2 to change the steering intensity and conduct a test on $MC^{2}$ for both OneVision-7B and Qwen2.5-VL-7B, reporting the $S_{\text{Vision}}$ or $S_{\text{Text}}$ scores in Table 10. We observe that performance drops for both models with decreased steering intensity, indicating insufficient steering. As intensity increases, the performance of OneVision-7B significantly drops, exhibiting clear over-steering at $\lambda=2.0$ which leads to destruction of language capabilities. Conversely, for Qwen2.5VL-7B, the steering effect continues to enhance up to $\lambda=1.75$, only significantly degrading beyond $\lambda=2.0$, demonstrating a wider safe steering margin in its representation space.
Table 10: The ablation study for the varied steering intensities.
Model 0.125 0.25 0.5 0.75 1.0 1.25 1.5 1.75 2.0 2.25 2.5 2.75 3.0
OneVision-7B 40.2 45.5 51.8 52.6 57.1 56.5 32.6 12.1 3.9 0.1 0.0 0.0 0.0
Qwen2.5VL-7B 40.8 44.1 49.4 53.2 63.6 72.3 76.0 70.7 45.8 15.2 16.2 12.5 7.8
<image> represents the visual input displayed below
User: $\langle\text{image}\rangle$ Text Context: The table was adorned with a vibrant bouquet of flowers and a charming ceramic sheep, while the surrounding chairs, crafted from smooth, polished wood, complemented the rustic yet elegant setting. In case of inconsistency, prioritize the text context.
Question: What are the chairs made of? A. wicker    B. wood
Assistant: $\langle\text{output}\rangle$
Vision-based answer: wicker
Text-based answer: wood
Baseline answer: A. wicker ✗
Proposed steering answer: B. wood ✓
Figure 12: Example of a cross-modal conflict sample from our diagnostic dataset. The highlighted text indicates the ground-truth modality specified by the instruction, which the proposed steering method successfully follows while the baseline fails.
Figure 13: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 23rd layer.
Figure 14: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 24th layer.
Figure 15: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 25th layer.
Figure 16: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 26th layer.
Figure 17: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 27th layer.
Figure 18: The attention scores of the generated token toward the vision and text contexts using InstDesign (Left) and the proposed method (Right) in the 28th layer.
E.4 Details for the application of the proposed method
E.4.1 Details of in-depth analysis for steering method
We provide the detailed description of the case for in-depth analysis for steering method in Figure 6 in Section 5.4. Besides, we also provide more attention analysis for the case in the different layers in Figure 1314151617 and 18. We observe that across all subsequent layers following steering modality preference towards text at the 21 th layer, our method significantly increases the model’s attention weight toward the text modality, surpassing the corresponding vision attention weight. This change clearly demonstrates that the steering mechanism successfully alters the modality preference by enhancing the model’s dependency on the text modality.
E.4.2 Latency and Memory
The proposed method consists of two phases, probing and steering. The probing phase is conducted offline and the resulting steering vector is cached for reuse. During the actual steering phase, we simply load this cached steering vector, which incurs minimal memory overhead and does not add meaningful computational cost to the inference process. We measure the single-sample inference latency (seconds) (without Flash-Attention acceleration and batch inference) for our method compared to the MLLM baseline (MLLM-only) in Table 11. The results show that the steering phase introduces negligible latency compared to the MLLM-only baseline, and the overhead is confined to the initial offline probing stage. Furthermore, all three methods require nearly identical memory requirements.
E.4.3 The prerequisites for implementation of the proposed method
Based on Representation Engineering (Greenblatt et al., 2023; Xu et al., 2024), the proposed method requires capturing an explicit modality preference direction vector to realize behavioral adjustment. The approach succeeds when such a vector can be reliably extracted, as seen in models like Qwen2.5VL-7B. However, the method fails in cases such as LLaVA-1.5-7B, primarily due to its limited ability to follow instructions for preference adjustment, which prevents the capture of a meaningful direction vector.
Table 11: The comparison of inference time between MLLM-only and the proposed method including probing and steering phases.
Dataset MC2 Phd-icc
MLLM-only 1.99 1.84
Probing (Offline) 2.21 2.12
Steering 2.00 1.84
Besides, we observe a localized performance drop in the Attribute subset of the Phd-icc benchmark in Table 3. We attribute this to the limitation of applying a single global steering vector, which may fail to accommodate instance-level granularity—where fine-grained, sample-specific features are required for optimal alignment. Despite these isolated cases, the overall benchmark performance improves, underscoring the effectiveness of our approach, especially considering that it requires no external data or fine-tuning.
E.5 The Further Experiment without modality preference prior
Our current approach intentionally select the steering direction based on known task requirements. This design has been proven to be pragmatic and effective for many real-world applications where the optimal modality is clear. In addition, our method can be readily integrated with a training-free priority detection method to enable dynamic preference selection. To demonstrate this, we conduct the following experiment:
Dataset Construction: We modify $MC^{2}$ dataset by degrading the quality of one modality context so that only one modality is reliable, and the ground-truth answer aligns with it. We use QA accuracy to measure model performance on this new dataset.
Task Design: 1) Each sample requires a specific reliable modality. 2) All samples share the same reliable modality in a task. Each task contains 200 samples.
Solution: We apply a causal analysis approach (Parcalabescu and Frank, 2024) to identify the reliable modality. For each sample, we first measure the change of predicted answer probability when removing either the image or the text context. The larger the drop, the more important that modality is for the given sample. For Task1, we determine the reliable modality for a specific sample by comparing the probability drops. For Task2, by aggregating the reliable modalities across all samples via majority voting, we determine the preferred modality for the specific task.
Results: For the identification of reliable modality, we achieve an accuracy of $85.3\%$ for all samples in Task1; we reach $100\%$ accuracy for task-level identification in Task2 (thus, performance on Task 2 is equivalent to knowing the steering preference in advance). Next, we evaluate the performance of the proposed method on Task1, measured by QA accuracy in Table 12.
The results show that steering with predicted preference yields significant gains over base models and closely matches the performance of the “preference prior” setting. This confirms that our method can be simply adapted to autonomously prioritize modalities based on input quality or task needs.
Table 12: The performance of the proposed method on the revised $MC^{2}$ without modality preference prior measured by Accuracy.
Method Task1
OneVision-only 25.4
+Steering with preference prior 40.7
+Steering with predicted preference 37.5
Qwen2.5VL-7B-only 49.1
+Steering with preference prior 62.7
+Steering with predicted preference 58.2